3. You see a gold ring in a flea market that you'd like to have, and the price seems right, so you buy it. On the drive bome you wonder if it's really gold. You do a quick experiment and find it takes 38.7 J of heat to raise the temperature of the 6.00 g ring from 2SC to 75°C. You know that the specific heat ofgold is 0.128Jg"С, is your ring really made of gold?


1
Expert's answer
2021-06-23T05:17:19-0400

Heat = mass × specific heat × change in temperature

38.7 = 6.00 × c × (75-25)

38.7 = 300 × c

c = 38.7/300

= 0.129J/g°C

The specific heat of ring calculated is slightly different than the specific heat of gold, so the ring might not be made up of gold.